Chapter 17, Problem 058 A sound source A and a reflecting surface B move directly toward each other. Relative to the air, the speed of source A is 28.9 m/s, the speed of surface B is 64.3 m/s, and the speed of sound is 326 m/s. The source emits waves at frequency 1230 Hz as measured in the source frame. In the reflector frame, what are (a) the frequency and (b) the wavelength of the arriving sound waves? In the source frame, what are (c) the frequency and (d) the wavelength of the sound waves reflected back to the source?
Answer #1

VA = 28.9 ms V= 326m/s UB= 64.3 m/s f = 1230Hz (a) From Doppler effect For reflector frame it 2 + ( + MB) $= 1230 ( 326 +64(C) for source frame. th a t ( VTVA) f = 1.62X103 / 326+28. 91 It = 2.19 x103 Hz a = 7/11 = 326 2.19x103 1x = 0.149m) 326-

Do let me know in case of any mistake..

I will rectify it asap..

Thank you .